OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 Numerical Answers Expected!

OpenStudy (anonymous):

b^2 − 4ac

OpenStudy (anonymous):

Now, 2x2 + 7x + 3 = 0 =>2x^2+6x+x+3=0 => 2x(x+3)+1(x+3)=0 =>(2x+1)(x+3)=0

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

so the answer is (2x+1)(x+3)=0

OpenStudy (anonymous):

Nope, Solve that for x

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

Either 2x+1=0 or x+3=0

OpenStudy (anonymous):

does that help?

OpenStudy (goformit100):

@rebecca1233 r u there>?